OpenStudy (anonymous):

15 + 102 • [12 – (12 – 3 • 2)] + 6

OpenStudy (misssmartiez):

back, still need assistance?

OpenStudy (anonymous):

yes please

OpenStudy (misssmartiez):

do 3 • 2 first ( mental math)

OpenStudy (anonymous):

6

OpenStudy (misssmartiez):

12 – 6.

OpenStudy (anonymous):

6

OpenStudy (f_jayyy):

Use PEMDAS...but since there are brackets do those first

OpenStudy (anonymous):

ya

OpenStudy (misssmartiez):

right, 15 + (102 •6) + 6

OpenStudy (anonymous):

no the parentheses are first

OpenStudy (misssmartiez):

do 102 x 6.

OpenStudy (f_jayyy):

Technically no...since the parentheses are inside the brackets you must do the brackets first

OpenStudy (seanstewart16):

you do parenthesis, and then brackets

OpenStudy (anonymous):

@MissSmartiez 612

OpenStudy (misssmartiez):

.. It really wouldn't count. YOU DO WHAT IS IN ( ) then subtract by 12 (which is in the brackets.

OpenStudy (f_jayyy):

yah ik but what i mean is you do what is inside the brackets

OpenStudy (anonymous):

ok

OpenStudy (misssmartiez):

612 + 15 then add six.

OpenStudy (misssmartiez):

That is what we did, F_Jay...

OpenStudy (anonymous):

633

OpenStudy (misssmartiez):

Good job!

OpenStudy (misssmartiez):

I believe that should be the answer, to the question, about 90% chance c:!

OpenStudy (anonymous):

:)

OpenStudy (misssmartiez):

You saw how we broke it down?

OpenStudy (anonymous):

yes

OpenStudy (misssmartiez):

If you still wish for my assistance, summon me.

OpenStudy (anonymous):

ok

OpenStudy (misssmartiez):

This is what you confuse, F_Jayyy. All the items we calculated was in the '[]' But we done parenthesis first because it was also inside the brackets, WITH PARENTHESIS. Then we subtracted the total with 12

OpenStudy (misssmartiez):

If you didn't realize, the item wouldn't change, because 12 - 6 is still 6.

OpenStudy (mathstudent55):

15 + 102 • [12 – (12 – 3 • 2)] + 6 = 15 + 102 • [12 – (12 – 6)] + 6 = 15 + 102 • [12 – 6] + 6 = 15 + 102 • 6 + 6 = 15 + 612 + 6 = 627 + 6 = 633